OpenStudy (anonymous):

1. A basketball player jumps off the floor during the tip-off. Show how to determine the amount of time that is required for the player to reach his maximum height if he leaves the floor with a velocity of 3.30 m/s. 2. A child throws a ball straight up into the air. If 4.2 seconds are required for the ball to go up and then come back to the child’s hands, how much time was required for the ball to reach its highest point?

OpenStudy (anonymous):

v = 3.30 m/s - (9.8 m/s^2) t v= 0 to get t child's ball y = v t - (1/2) a t^2 = 0 solutions are t=0 [start] v = (1/2) a t t = 2 v/a the total time is twice the time (v/a) it takes for the upward velocity to reach 0. Thus....

OpenStudy (anonymous):

Thank you, folks.